Lisa consumes only pizzas and burritos. In equilibrium, her marginal utility of pizza is 28 and her marginal utility of a burrito is 21. The price of a pizza is $3. Instructions: Round your answer to 2 decimal places. The price of a burrito is ___?
Answer #1

Ans) $ 2.25

At equilibrium, Marginal utility of pizza / marginal utility of burrito = price of pizza / price of burrito

28/21 = 3 / price of burrito

Price of burrito = 21*3/28 = 2.25